OpenStudy (bossimbacon):

Which ordered pair is the solution to the system of equations? 2x+y=-8 x-y=-4 A. (−3, −2) B. (−1, 3) C. (1, 5) D. (−4, 0)

OpenStudy (bossimbacon):

@SolomonZelman

OpenStudy (anonymous):

solving a problem like this is usually a matter of solving for one variable and plugging that into the other equation.

OpenStudy (anonymous):

Step 1) solve for a variable x-y=-4 x=-4+y Step 2) substitute into second equation and solve for the other variable. 2x+y=-8 2(-4+y)+y=-8 -8+2y+y=-8 3y=0 ---> y=0 Step 3) plug in solved variable to find unknown variable y=0 x-y=-4 --->x-0=-4 x=-4 Therefore (-4,0)

OpenStudy (bossimbacon):

wow thank you! can you help me with one more?

OpenStudy (anonymous):

sure but do you understand what I did?

OpenStudy (bossimbacon):

yeah i do:D

OpenStudy (bossimbacon):

you listed it out for me I'm dyslexic so that really helped a lot

OpenStudy (anonymous):

ok cool. whats up?

OpenStudy (bossimbacon):

Which ordered pair is the solution to the system of equations? 2x-5y-5 x+2=11 A. (3, 4) B. (5, 3) C. (7, 2) D. (15, 7)

OpenStudy (anonymous):

well, you tell me what is the first thing you would do here?